# FloorplanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v2

@@ Abstract
We introduce FloorplanQA, a diagnostic benchmark for evaluating spatial reasoning in large-language models (LLMs). FloorplanQA is grounded in structured representations of indoor scenes, such as (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts. The benchmark covers core spatial tasks, including distance measurement, visibility, path finding, and object placement within constrained spaces. Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed in shallow queries, they often fail to respect physical constraints, preserve spatial coherence, though they remain mostly robust to small spatial perturbations. FloorplanQA uncovers a blind spot in today’s LLMs: inconsistent reasoning about indoor layouts. We hope this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
@@ 1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over plausible, physically feasible environments such as indoor layouts remains poorly understood. In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce FloorplanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floorplans represented in structured text-based formats. Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. FloorplanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. FloorplanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding, in order to measure their unassisted capabilities. This baseline is important because even in tool-rich systems, models benefit from some unaided spatial ability to anticipate outputs and avoid trivial errors.
@@
  • We introduce a dataset of 2,000 structured 2D layouts, including 600 each from synthetically generated kitchens, living rooms, and bedrooms, plus 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023), providing a realism check. All are represented in JSON and paired with spatial reasoning questions.
@@ 2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR Johnson et al. (2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations. In real-world settings, SpatialSense Yang et al. (2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI Chevalier-Boisvert et al. (2019), ALFRED Shridhar et al. (2020), and Room-to-Room (R2R) Anderson et al. (2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA Azuma et al. (2022) and 3DSRBench Ma et al. (2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
@@
Vision-language models have advanced spatial reasoning but often handle it qualitatively. The VQA dataset Antol et al. (2015) challenges models to answer questions about images, while VL-T5 Cho et al. (2021) unifies vision-and-language tasks via text generation. Recent work on 3D scene graphs Armeni et al. (2019) introduces structured representations of environments, facilitating spatial reasoning. However, these approaches may miss fine-grained geometric details necessary for precise spatial inference. Efforts like SpatialVLM Chen et al. (2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
@@
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT Feng et al. (2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al. (2024) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome Fu et al. (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but often produces implausible object placement due to non-convergent simulated annealing. LayoutVLM Sun et al. (2024) and FirePlace Huang et al. (2025) improve layout generation via optimization and constraint solving, respectively. But they assess output realism, not the model’s ability to infer constraints directly. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
@@
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al. (2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks. Additionally, benchmarks such as BALROG Paglieri et al. (2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence. Recent 3D-LLM surveys such as (Ma et al., 2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning. FloorplanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
@@
FloorplanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts. Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, FloorplanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
@@ 3 Method
3.1 Synthetic Layout Generation
Our initial aim was to use publicly available real-world floorplan datasets. However, a comprehensive review revealed significant limitations, as several prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse. Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Google Gemini Robotics Team, 2025). Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: the data-generation step is one-off and separate from evaluation. For evaluation, models solely produce answers to the benchmark questions, and correctness is computed by deterministic geometric routines against ground-truth solutions derived from the layouts, so no model output is used to generate data or to grade itself.
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning. These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints. The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table; see Appendix B.3 for the full set of cases. Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
3.2 Layout Extraction from HSSD Dataset
To complement the synthetically generated layouts, we further incorporated 200 layouts extracted from the Habitat Synthetic Scenes Dataset (HSSD-200). HSSD provides 211 high-quality, human-authored 3D scenes designed with the Floorplanner interface and populated with 18,656 objects across 466 semantic categories. Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts.
For our purposes, we project each 3D scene to a 2D floorplan and retain only select structural and furniture elements. Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details. We then use an $\alpha$-convex hull (Asaeedi et al., 2014; Edelsbrunner & Mücke, 1994) to smooth object boundaries, yielding polygonal layouts that are not restricted to axis-aligned rectangles. This step is necessary because raw HSSD projections often produce overly dense polygons, with redundant vertices along straight or nearly straight segments; applying an $\alpha$-hull reduces spurious complexity while preserving concavity, which avoids unnecessary token overhead in downstream LLM processing. This ensures compatibility with our synthetic layouts, while maintaining the richer geometric variety of HSSD.
3.3 Unified Dataset
Together, these two sources yield a dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD. The two subsets share a unified polygonal representation, enabling consistent downstream processing. Figure 1 shows examples from both sources while Table 1 summarizes room, style, and geometry distributions across the synthetic subset with the figure next to it providing a breakdown by room type.
Table 1: Layout distribution in FloorplanQA by room type, style, and geometry, with room-wise totals.
@@ 3.3 Unified Dataset
Figure 1: Representative layouts from FloorplanQA. Generated: kitchen, living room, bedroom. HSSD: last image. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
@@ 3.4 Question Taxonomy and Prompting
To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line. For example, a distance query is phrased as:
Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR. Otherwise, responses must follow the scheme, for example: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.
@@
Repositioning Calculate how far the ottoman be moved in the left direction until it touches another object or the wall N Dynamic
@@
Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from centroid of the stove and ending at the centroid of door S Dynamic
@@
Layouts are represented in a structured JSON format. Each entry contains a layout_id, the room_type, and explicit geometric descriptions. The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments. Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list. All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon. In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation. Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward). The global origin $(0,0)$ is not fixed to the layout’s lower-left corner and may vary across layouts. The prompt used to generate examples according to this schema is in Appendix H. We group these questions into three reasoning categories. Metric tasks require explicit numerical computation, such as calculating centroids, measuring distances between objects, or evaluating the angle between an inter-object vector and a reference axis. Topology category involves geometric and relational reasoning, including checking placement feasibility, computing free space, or identifying whether an object blocks the direct line of sight between two others. Dynamic category addresses layout-changing procedures, such as repositioning an object until contact with a boundary or another object, or computing a valid collision-free path between two objects. These categories are intended to capture the core modes of spatial reasoning in FloorplanQA, ranging from low-level geometric calculation to higher-level relational and procedural inference. While not strictly disjoint, they provide a diagnostic framework for analyzing model behavior and diagnosing failure modes. In addition to categorizing by reasoning type, each task is also associated with an answer format code that specifies the expected output structure and the corresponding scoring rule. Scalar outputs (N) are scored by relative error with a default tolerance of 2%; for complex area-computation tasks (e.g., Free Space), the tolerance is relaxed to 5%. Sequence outputs (S) are evaluated with a Fréchet threshold of 0.6 m, approximating minimal human clearance, and must be valid (collision-free, no overlaps). List outputs (L) are evaluated by set equality. Together, the categories and format codes define the taxonomy summarized in Table 2, which reports task coverage and examples for each case.

### v3

@@ Abstract
We introduce FloorplanQA, a diagnostic benchmark for evaluating spatial reasoning in large-language models (LLMs). FloorplanQA is grounded in structured representations of indoor scenes (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts. The benchmark covers core spatial tasks, including distance measurement, visibility, path finding, and object placement within constrained spaces. Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed on shallow queries, they often fail to respect physical constraints and preserve spatial coherence, though they remain mostly robust to small spatial perturbations. FloorplanQA uncovers a blind spot in today’s LLMs: inconsistent reasoning about indoor layouts. We hope this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
Spatial Reasoning, Layout Reasoning, Scene Understanding, Structured Scene Representations, Benchmark, Large Language Models (LLMs)
@@ 1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over plausible, physically feasible environments such as indoor layouts remains poorly understood. In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce FloorplanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floorplans represented in structured text-based formats. Each instance consists of a symbolic (JSON-encoded) layout paired with natural language questions that require the model to reason about spatial and geometric aspects of the scene, compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. FloorplanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments. In contrast to many layout benchmarks that focus on qualitative relations or output realism, FloorplanQA emphasizes geometric consistency under explicit metric, clearance, and collision constraints.
Although LLMs can increasingly be used in tool-assisted pipelines, for example, to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. FloorplanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding, in order to measure their unassisted capabilities. This baseline is important because even in tool-rich systems, models benefit from some unaided spatial ability to anticipate outputs and avoid trivial errors.
@@
  • We introduce a dataset of 2,000 structured 2D layouts, including 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023) based on real floorplans, and 600 each from synthetically generated kitchens, living rooms, and bedrooms. All are represented in JSON and paired with spatial reasoning questions.
@@ 2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR (Johnson et al., 2016) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations. In real-world settings, SpatialSense (Yang et al., 2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI (Chevalier-Boisvert et al., 2018), ALFRED (Shridhar et al., 2019), and Room-to-Room (R2R) (Anderson et al., 2017) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA (Azuma et al., 2021) and 3DSRBench (Ma et al., 2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
@@
Vision-language models have advanced spatial reasoning but often handle it qualitatively. The VQA dataset (Agrawal et al., 2015) challenges models to answer questions about images, while VL-T5 (Cho et al., 2021) unifies vision-and-language tasks via text generation. Recent work on 3D scene graphs (Armeni et al., 2019) introduces structured representations of environments, facilitating spatial reasoning. However, these approaches may miss fine-grained geometric details necessary for precise spatial inference. Efforts like SpatialVLM (Chen et al., 2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
@@
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT (Feng et al., 2023) leverages large language models for compositional visual planning and layout generation, while Holodeck (Yang et al., 2023) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome (Wen et al., 2023) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but often produces implausible object placement due to non-convergent simulated annealing. LayoutVLM (Sun et al., 2024) and FirePlace (Huang et al., 2025) improve layout generation via optimization and constraint solving, respectively. But they assess output realism, not the model’s ability to infer constraints directly. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
@@
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models (Yamada et al., 2023), which assesses spatial reasoning through natural language descriptions of qualitative relations (e.g., ”left of,” ”above”), without numeric coordinates or metric computation. Additionally, benchmarks such as BALROG (Paglieri et al., 2025) test agentic reasoning in game environments (NetHack, BabyAI, etc.), where spatial reasoning emerges implicitly through sequential navigation commands rather than explicit geometric inference. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning over structured coordinate-based layouts, requiring models to compute distances, angles, collision-free paths, and spatial unions directly from symbolic representations. Recent 3D-LLM surveys, such as Ma et al. (2024b), cover tasks like embodied navigation and interaction, but not symbolic geometric reasoning from structured layouts.
@@
FloorplanQA addresses this gap by providing explicit spatial representations (object coordinates and dimensions). Unlike prior benchmarks relying on raw images or commonsense spatial language, it tests precise geometric tasks: computing distances, angles, collision-free paths, and spatial unions directly from symbolic input.
@@ 3 Method
Figure 1: (Left) The illustration of the overall breakdown by room type for the entire 2,000-layout benchmark, encompassing both the synthetic component and the layouts extracted from the HSSD. (Right) Detailed distribution of the synthetic subset of the FloorplanQA benchmark (1,800 layouts) across room type, internal style, and geometric configuration.
3.1 Layouts Used in This Study
We aim to create a benchmark based on floorplan datasets, with some critical requirements; first, they need to be publicly accessible and second, they need to include furnishings (couches, tables, chairs, etc) at a minimally plausible level. These requirements matter because FloorplanQA’s questions require object-level geometry (sizes, orientations, clearances), and we need a dataset that others can actually obtain, so that results are reproducible. In practice, we were unable to find a dataset that meets these requirements, as several prominent datasets such as SUNCG (Song et al., 2016) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2019), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse. Given these legal, practical, and methodological constraints, we found a combination of data derived from HSSD Khanna et al. (2023) based on real data, extended with LLM-based synthetic data generation, to be the most viable alternative.
We incorporated 200 layouts extracted from HSSD (Khanna et al., 2023). HSSD provides 211 high-quality, human-authored 3D scenes populated with 18,656 objects across 466 semantic categories. Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts. Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details. The original data is 3D, so we project each 3D scene to a 2D floor plan. This can result in dense polygons for each asset (repeated points, points on straight lines, excessive tokens), which would confound our ability to test 2D geometric reasoning, so we use the Douglas-Peucker algorithm (Douglas and Peucker, 1973) with $\epsilon=0.01$m to simplify the shapes (see Figure 4 for an example, Appendix C).
To increase the diversity of our benchmark, we generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro (at the time the most capable publicly available variant), with all layouts generated up-front and fixed prior to evaluation. The Gemini 2.5 Pro model has been trained on spatial reasoning and robotic manipulation tasks (Google Gemini Robotics Team, 2025), making it well-suited for generating geometrically plausible indoor layouts. Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: (1) layout generation is a one-time process separate from benchmark evaluation, (2) all ground-truth answers are computed deterministically using geometric algorithms (not model outputs), and (3) correctness is verified via rule-based spatial calculations.
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to clearance, symmetry, and zoning (see Appendix B for more details). These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints. The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table. Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
3.2 Unified Dataset
@@
Figure 2: Representative layouts from FloorplanQA. HSSD: living room (first image) Generated: bedroom, kitchen and living room. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
Together, these two sources yield a publicly available dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD. The two subsets share a unified polygonal representation, enabling consistent downstream processing. Figure 2 shows examples from both sources, while Figure 1 (left) plots the room-type distribution; Figure 1 (right) reports the corresponding room/style/geometry counts for the synthetic subset.
Layouts are represented in a structured JSON format. Each entry contains a layout_id, the room_type, and explicit geometric descriptions. The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments. Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list. All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon. In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation. Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward), with origin at $(0,0)$. During synthetic data generation, layouts are initially produced in a top-left origin frame (+y downward) following Gemini 2.5 Pro’s native output format. All generated layouts are automatically transformed to the right-handed coordinate system before benchmark construction, ensuring consistency across synthetic and HSSD data. HSSD layouts are similarly transformed from their original 3D projection coordinates to this unified 2D frame. The prompt used to generate examples according to this schema is in Appendix J.
@@ 3.3 Question Taxonomy and Prompting
Repositioning Calculate how far the ottoman can be moved in the left direction until it touches another object or the wall N Dynamic
@@
Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from the centroid of the stove and ending at the centroid of the door S Dynamic
@@
We group these questions into three reasoning categories. Metric tasks require explicit numerical computation, such as calculating centroids, measuring distances between objects, or evaluating the angle between an inter-object vector and a reference axis. Topology category involves geometric and relational reasoning, including checking placement feasibility, computing free space, or identifying whether an object blocks the direct line of sight between two others. Dynamic category addresses layout-changing procedures, such as repositioning an object until contact with a boundary or another object, or computing a valid collision-free path between two objects.
These categories are intended to capture the core modes of spatial reasoning in FloorplanQA, ranging from low-level geometric calculation to higher-level relational and procedural inference. While not strictly disjoint, they provide a diagnostic framework for analyzing model behavior and diagnosing failure modes.
In addition to categorizing by reasoning type, each task is also associated with an answer format code that specifies the expected output structure and the corresponding scoring rule. Scalar outputs (N) are scored by relative error with a default tolerance of 2%; for complex area-computation tasks (e.g., Free Space), the tolerance is relaxed to 5%. These tolerances are chosen to accommodate minor numerical instability in LLM outputs while remaining strict enough to distinguish correct spatial reasoning from approximation. Sensitivity studies supporting these thresholds, including tolerance sweeps showing that they affect only absolute accuracies and not model rankings, are provided in Appendix G.1. List outputs (L) are evaluated by set equality. Sequence outputs (S) must be valid (collision-free with required clearance) and are additionally scored by Fréchet distance to the reference path. We use a threshold of 0.6 m, which tolerates minor deviations in waypoint placement while rejecting qualitatively different routes (e.g., opposite sides of an obstacle). Threshold sweeps are reported in Appendix G.2.
To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line. Some models expose API-level mechanisms for structured output enforcement. We instead encode the same schema as formatting requirements in the prompt for all models to keep the prompting interface fixed across systems. For example, a distance query is phrased as:
Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR. Otherwise, responses must follow the scheme, for example: Prompt: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.